OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $2.10. If the number of dimes was tripled and the number of nickels was increased by 24, the value of the coins would be $5.70. How many nickels and dimes does he have?

OpenStudy (cherylim23):

So we know that $2.10 = 210c and $5.70 = 570c Let the number of nickels be x and number of dimes be y. At first, you had x nickels and y dimes, worth 210c Hence 5x + 10y=210 --(1) Later you had x+24 nickels and 3y dimes, worth 570c. Hence 5(x+24) + 10 (3y) = 570 5x + 120 + 30y = 570 5x+30y = 450 --(2) Solve equations (1) and (2) simultaneously (I recommend by elimination), you should get x=18 and y=12. You should know your answer by now.

OpenStudy (anonymous):

yes that helped, thank you!